OpenStudy (anonymous):

Solve: 8x2 + 14x = 15

OpenStudy (ashleyisakitty):

x=-5/2, x= 3/4

OpenStudy (anonymous):

8x2 = 16. 14x (Do the inverse operation which would be division) 14/14 (They cancel out so just forget about 14. 15/14 = (and that is your answer)